The hundreds digit of a 3-digit number is 2 less than the tens digit, and the ones digit is 2 less than the hundreds digit. If twice the sum of the digits decreased by 10 is 20, what is the number?
Answer by jorel1380(3719)

Let n be the tens digit of the number. Then the hundred digit is n-2, and the ones digit is n-2-2, or n-4. So:
2(n+n-2+n-4)-10=20
6n-22=20
6n=42
n=7
n-2=5
n-4=3
So the number is 573. ☺☺☺☺